OpenStudy (anonymous):

there's this one vertical motion word problem: A falling object encounters air resistance that is proportional to its velocity. The acceleration due to gravity is -9.8 meters per second per second. The net change in velocity is dy/dt=kv-9.8. (a) find the velocity of the object as a function of time if the intial velocity is v(subscript "0"). (b)Use the result of part (a) to find the limit of the velocity as t approaches infinity. (c) Intergrate the velocity function found in part (a) to find the position function "s".

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large \frac{dv}{dt}=kv-9.8\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

so, I would integrate the dy/dt=kv-9.8 to get the velocity equation for a...so, v(subscript 0)= kv^2 -9.8 v? because the K is just a constant, even if divided by 2, wouldn't really show up... and yesh, exactly like that ^^

OpenStudy (anonymous):

You will get an exponential approach to a terminal velocity v = (vfinal)(1-exp(-kt))

ganeshie8 (ganeshie8):

Alright, you're mixing y's and v's

ganeshie8 (ganeshie8):

so, there are NO y's in the question anywhere right ?

ganeshie8 (ganeshie8):

its all about v's

ganeshie8 (ganeshie8):

?

OpenStudy (anonymous):

oh! yeah~ hehe whoops many apologies~

ganeshie8 (ganeshie8):

\(\large \frac{dv}{dt}=kv-9.8 \)

ganeshie8 (ganeshie8):

:) just making sure we'rent missing anything... thats all :)

ganeshie8 (ganeshie8):

okay, Velocity and Time are variables here

ganeshie8 (ganeshie8):

separate variables

ganeshie8 (ganeshie8):

\(\large \frac{dv}{dt}=kv-9.8 \) \(\large \frac{dv}{kv-9.8}= dt \)

ganeshie8 (ganeshie8):

Integrate

ganeshie8 (ganeshie8):

\(\large \int \frac{dv}{kv-9.8}= \int dt \)

ganeshie8 (ganeshie8):

and you're right, k is just a constant

OpenStudy (anonymous):

oh so then the ln rule thing, right?

ganeshie8 (ganeshie8):

yup!

OpenStudy (anonymous):

and so,....oh...hold on... so then where does the k go??? wud it go outside the ln like this? kln|kv-9.8|+c?

ganeshie8 (ganeshie8):

differentiate kln|kv-9.8|+c, and check if u get back ur 1/(kv-9.8)

ganeshie8 (ganeshie8):

if u get back, then ur integration is correct, otherwise - think again :)

OpenStudy (anonymous):

oh, uhm, would the ln be divided by k?

ganeshie8 (ganeshie8):

you got it :)

OpenStudy (anonymous):

ok so then the answer to part a would be ln|kv(subscript 0)-9.8|/k +c =V

ganeshie8 (ganeshie8):

\(\large \frac{dv}{dt}=kv-9.8\) \(\large \frac{dv}{kv-9.8}= dt\) \(\large \int \frac{dv}{kv-9.8}= \int dt \) \(\large \frac{\ln |kv-9.8|}{k}= t + c \) \(\large \ln |kv-9.8|= kt + c \) \(\large kv-9.8= e^{kt + c} \) \(\large kv-9.8= ce^{kt} \)

ganeshie8 (ganeshie8):

solve \(\large v\)

ganeshie8 (ganeshie8):

and see that, i have change the place of \(c\) watever way i felt good

OpenStudy (anonymous):

ohh ok~ so then\[v=\frac{ ce^{kt}+9.8 }{ k }\]

ganeshie8 (ganeshie8):

Correct, lets find the value of \(c\) now :)

ganeshie8 (ganeshie8):

read part a) question again we're given some initial value right ?

ganeshie8 (ganeshie8):

(a) find the velocity of the object as a function of time if the intial velocity is v(subscript "0").

ganeshie8 (ganeshie8):

initial velocity = \(V_0\) wat does that mean ?

OpenStudy (anonymous):

is that supposed to be given?

ganeshie8 (ganeshie8):

that means, when time, \(t = 0\), the velocity of object \(v = V_0\)

OpenStudy (anonymous):

oh, rght~XD

ganeshie8 (ganeshie8):

so, \((0, V_0)\) is a point on out velocity function \(\large v\)

ganeshie8 (ganeshie8):

plugit in and find the value of \(c\)

ganeshie8 (ganeshie8):

\(\large \frac{dv}{dt}=kv-9.8\) \(\large \frac{dv}{kv-9.8}= dt\) \(\large \int \frac{dv}{kv-9.8}= \int dt \) \(\large \frac{\ln |kv-9.8|}{k}= t + c\) \(\large kv-9.8= e^{kt + c}\) \(\large kv-9.8= ce^{kt}\) \(\large v= \frac{ce^{kt}+9.8}{k}\) plugin \((0, V_0)\) \(\large V_0 = \frac{ce^{0t}+9.8}{k} \implies c = V_0k - 9.8\) So, the particular velocity function is : \(\large v= \frac{(V_0k-9.8)e^{kt}+9.8}{k}\)

ganeshie8 (ganeshie8):

we're done with part a)

ganeshie8 (ganeshie8):

let me knw if smthng doesnt make sense...

OpenStudy (anonymous):

ohok got it XP haha redid some stuff to make sure i got it ^^

ganeshie8 (ganeshie8):

good :)

ganeshie8 (ganeshie8):

before moving to part b) let me ask u a simple question

OpenStudy (anonymous):

ok go for it ^^

ganeshie8 (ganeshie8):

lets assume climbing up is positive, here is the question : when you're FALLING DOWN, whats the "sign" of ur velocity ?

OpenStudy (anonymous):

OH negaitve!

ganeshie8 (ganeshie8):

good, now look at our final velocity equation : \(\large v= \frac{(V_0k-9.8)e^{kt}+9.8}{k}\)

ganeshie8 (ganeshie8):

since the object is always falling down, can we say \(\large v\) needs to be negative always ?

OpenStudy (anonymous):

yes~ and so would we have to make k a negative number?

ganeshie8 (ganeshie8):

you got it !!

ganeshie8 (ganeshie8):

\(k\) is a negative number, having that knowledge we're completely ready to pursue part b)

ganeshie8 (ganeshie8):

(b)Use the result of part (a) to find the limit of the velocity as t approaches infinity.

ganeshie8 (ganeshie8):

\(\large \lim \limits_{t\to \infty }v= \lim \limits_{t\to \infty }\frac{(V_0k-9.8)e^{kt}+9.8}{k}\)

OpenStudy (anonymous):

the exponent of e gets infintely negative, and so that would mean that Vk-9.8 is etting huge,

ganeshie8 (ganeshie8):

what do u get ?

OpenStudy (anonymous):

er, small, srry

ganeshie8 (ganeshie8):

what is the value of \(e^{-\infty}\) ?

ganeshie8 (ganeshie8):

i meant, what value it approaches..

OpenStudy (anonymous):

1/something super big, so it's approaching 0?

OpenStudy (anonymous):

ohhhhh hehe whoops omg i need sleep after this problem~

ganeshie8 (ganeshie8):

\(\large \lim \limits_{t\to \infty }v= \lim \limits_{t\to \infty }\frac{(V_0k-9.8)e^{kt}+9.8}{k}\) \(\large ~~~~~~~~~= \frac{(V_0k-9.8)(0) +9.8}{k}\) \(\large ~~~~~~~~~= \frac{9.8}{k}\)

OpenStudy (anonymous):

oh wait yeah I was right, actually ^^

ganeshie8 (ganeshie8):

yes you're right :)

ganeshie8 (ganeshie8):

btw, its an amazing result, if you haven't thought of it before

ganeshie8 (ganeshie8):

9.8/k is a constant ! the free falling body's velocity will not continue to increase forever, it reaches a limit and stays fixed there

OpenStudy (anonymous):

wattttttt!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! so there's a natural limit to how fast we can fall?

OpenStudy (anonymous):

that is interesting indeed...uhm... so onto part c, or?

ganeshie8 (ganeshie8):

Suppose ur plane crashed and you're falling down freely; then your velocity will increase in the start... but after some point it approaches a fixed value : air pushes u up gravity pulls u down how much force air exerts on u depends on ur velocity, so after u reach certain speed, the air and gravity effects balance each other out

ganeshie8 (ganeshie8):

(c) Intergrate the velocity function found in part (a) to find the position function "s".

ganeshie8 (ganeshie8):

\(\large v= \frac{(V_0k-9.8)e^{kt}+9.8}{k}\) Integrating velocity gives u position \(\large \int v dt= \int \frac{(V_0k-9.8)e^{kt}+9.8}{k} dt\) \(\large s = \frac{1}{k} \left( \int (V_0k-9.8)e^{kt}+9.8 ~dt \right)\)

ganeshie8 (ganeshie8):

piece of cake for u

OpenStudy (anonymous):

is it (1/k) (vk^2 e^(kt) - 9.8 ke^kt + 9.8)?

OpenStudy (anonymous):

whoops, last one is 9.8t

OpenStudy (anonymous):

NONO I did that WRONG

OpenStudy (anonymous):

it's actually v(sub 0 ) e^kt -9.8 e^kt + 9.8t all divided by k

OpenStudy (anonymous):

middle term is divided by k

ganeshie8 (ganeshie8):

\(\large s = \frac{1}{k} \left( \int (V_0k-9.8)e^{kt}+9.8 ~dt \right) \) \(\large s = \frac{1}{k} \left( (V_0k-9.8)\frac{e^{kt}}{k}+9.8t + c \right) \)

OpenStudy (anonymous):

yesh~ eexcept I forgot the c...hehe

ganeshie8 (ganeshie8):

^^constant comes up... cuz we dont knw from where we have started falling down... we're not given any initial position... so \(c\) sticks there

OpenStudy (anonymous):

ohh ok ^^ thank you lots!!! oh hey btw b4 we part, I have a random question~

ganeshie8 (ganeshie8):

shoot

OpenStudy (anonymous):

Have you lived in Alabama(US) before?

ganeshie8 (ganeshie8):

nooo... i never moved out of india till now why do u ask ha ?

OpenStudy (anonymous):

ohhh hahah you sounded like one of my teachers that I knew quite a while back ^^ woah india? You'e in india?

OpenStudy (anonymous):

is that where that "ha" comes from?

ganeshie8 (ganeshie8):

lol xD yes

OpenStudy (anonymous):

just curio~ XD

ganeshie8 (ganeshie8):

you're in Alabama now is it... must be very late over there lol u should sleep now... otherwise u wont be fresh for exam tomorrow

OpenStudy (anonymous):

^^ actually I'm in New Mexico hehe :D ok then~ thank you~ I'll cya around XP

ganeshie8 (ganeshie8):

have good sleep :) good luck for exam !! bye byeeeeee

OpenStudy (anonymous):

^^ I probs wont go to slp until later because I have to rememorize the all of the trig and inverse trig identities for deriv and integrals~ sigh :) thnx have a good afternoon XP

ganeshie8 (ganeshie8):

okay you're a machine lol !! i can give a hand.. when u need me just holler... il be around :)

OpenStudy (anonymous):

awesome ^^